Find an identity for cos(4t) in terms of cos(t).
cos(4t) =

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

look up Half-angle formulas on Wikipedia.com

Answer by ikleyn(53925)

.

Use the general trigonometric formula


    cos(2a) = 2cos^2(a) - 1.


twice.  It gives 


    cos(4t) = 2cos^2(2t) - 1 = 2*(2cos^2(t)-1)^2 - 1 = 2*(4cos^4(t) - 4cos^2(t) +1) - 1 = 

            = 8cos^4(t) - 8cos^2(t) + 1.     ANSWER

Solved and carefully explained.